OpenStudy (anonymous):

Find the largest number obtainable as the product of positive integers whose sum is 1976.

Parth (parthkohli):

Is this a brilliant.org question?

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

is it a problem to ask brilliant.org questions on OS

Parth (parthkohli):

I guess it'd be \(988 \times 988\).

OpenStudy (anonymous):

let \[x_1,x_2,....,x_n \space be \space the \space integers \] \[AM \ge GM\] \[\huge \frac{x_1+x_2+...+x_n}{n} \ge \sqrt[n]{x_1x_2...x_n}\] \[\huge \frac{1976^n}{n} \ge x_1x_2...x_n\]

Parth (parthkohli):

When you want to find the highest such number, just halve the sum.

OpenStudy (anonymous):

denominator n^n

Parth (parthkohli):

Oh myy...

OpenStudy (anonymous):

wats the problem

Parth (parthkohli):

I am not too good at all those applications . . . I just know that the numbers are \(988\) and \(988\)

OpenStudy (anonymous):

okay i will post the solution but its so much of analysis

OpenStudy (anonymous):